# spacing=nonfrench

### v1

@@ 6.1 Rogue Extension Negotiation Attack
AsyncSSH clients support the server-sig-algs and global-requests-ok extensions. Hence, the attacker can try to downgrade the algorithm used for client authentication by meddling with the value of server-sig-algs.
@@ 6.2 Rogue Session Attack
Figure 5: Rogue Session Attack on AsyncSSH: The MitM injects a malicious authentication request before the handshake completes. The client-side extension information message is deleted to account for the change in sequence numbers. By delaying the authentication request sent by the client, the MitM ensures that the malicious one is being processed. Any additional authentication requests are silently ignored.
The rogue session attack targets any SSH client connecting to an AsyncSSH server, on which the attacker must have a shell account. The goal of the attack is to log the client into the attacker’s account without the client being able to detect this. At that point, due to how SSH sessions interact with shell environments, the attacker has complete control over the remote end of the SSH session. The attacker receives all keyboard input by the user, completely controls the terminal output of the user’s session, can send and receive data to/from forwarded network ports, and is able to create signatures with a forwarded SSH Agent, if any. The result is a complete break of the confidentiality and integrity of the secure channel, providing a strong vector for a targeted phishing campaign against the user. For example, the attacker can display a password prompt and wait for the user to enter the password, elevating the attacker’s position to a MitM at the application layer and enabling impersonation attacks.
The messages exchanged during the attack are depicted in Figure 5. The attacks work by the attacker injecting a chosen UserAuthRequest before the client’s NewKeys. The UserAuthRequest sent by the attacker must be a valid authentication request containing his credentials. The attacker can use any authentication mechanism that does not require exchanging additional messages between client and server, such as password or publickey. Due to a state machine flaw, the AsyncSSH server accepts the unauthenticated UserAuthRequest message and defers it until the client has requested the authentication protocol.
To avoid a race condition between the UserAuthRequest sent by the client and UserAuthRequest injected by the attacker, the attacker delays the client’s UserAuthRequest until after the server signaled a successful authentication in response to the injected UserAuthRequest. The AsyncSSH server silently ignores any additional authentication request after a successful authentication.
@@
  2. 2.
    Suppose the server sends an extra message before ServiceAccept. In that case, the attacker can delete that message from the channel after injecting an additional Unknown message to the client before NewKeys, triggering a Unimplemented response. This increases both $\mathsf{C.Snd}$ and $\mathsf{C.Rcv}$, moving the send count deficit from the client to the server.
@@
We successfully evaluated the attack against AsyncSSH 2.13.2 as a server, connecting to with AsyncSSH 2.13.2 and OpenSSH 9.4p1.
@@ 7 SSH Deployment Statistics
To estimate the impact of the prefix truncation attacks, we scan for the SSH servers preferring or supporting the vulnerable encryption modes. Similarly, to estimate the impact of the extension downgrade attack, we scan for servers sending ExtInfo messages.
@@
For scanning, we used ZMap [14] and ZGrab2 [13] on port 22 of the entire IPv4 address space. The first scan was performed over two days in early October 2023, totaling 15.164M SSH servers.
As ZGrab2 cannot capture SSH extensions, we performed a second scan at the end of June 2023 using a custom tool on a subset of $2^{20}$ open ports discovered by ZMap. The scan covered a total of 830k servers. All data relating to the use of extension negotiation in SSH is sourced from this scan.
@@
In Table 1, we show the number of servers preferring and supporting various encryption modes. A cipher is preferred if placed first in the list of supported algorithms.
We find that, by far, the most preferred encryption cipher is ChaCha20-Poly1305, with 57.64% listing this algorithm first. This is followed by AES-CTR (31.53%) and, with some distance, by AES-GCM (8.04%) and AES-CBC (1.56%).
@@
As non-AEAD ciphers must be combined with a MAC, we also evaluate which authenticated encryption modes are preferred and supported by the servers. The numbers for the AEAD modes ChaCha20-Poly1305 (57.64%) and GCM (8.04%) are identical to those for encryption modes, as the MAC is already integrated. Preference for CTR modes is split into a majority for CTR-EaM (26.14%) and a minority for CTR-EtM (5.46%). Preference for CBC modes is mostly CBC-EaM (2.37%), while only a marginal share of servers prefers CBC-EtM (0.09%).
In summary, 57.73% of all servers prefer an authenticated encryption mode vulnerable to our attacks.
Looking at the support for authenticated encryption modes vulnerable to our attacks, we find that 67.58% of all servers support ChaCha20-Poly1305, while 17.24% support CBC-EtM. In total, 77% support at least one vulnerable mode, and 7% support both (these numbers are not shown in the table).
Cipher Mode and MAC Preferred Supported
@@
Other 4k 0.28% - -
Unknown / No KexInit 34k 0.23% - -
@@
We also looked at SSH extensions offered by servers before user authentication; see Table 3. We can see that 76.81% of all servers send the server-sig-algs extensions to indicate support for better signature schemes for client public key authentication. Furthermore, 8.8% send the publickey-hostbound extension, improving security for multi-hop authentication authentication using SSH agent. Both extensions provide opportunities for downgrade attacks, as their absence can weaken the strength of the authentication.
@@
Table 3: SSH extensions offered by servers after the initial handshake, @openssh.com abbreviated to @. Extensions sent by clients, and by servers upon successful client authentication are not included.
@@ 8 Suggested Countermeasures
As a stop-gap measure, the vulnerable cipher modes can be disabled. Suitable alternatives are AES-GCM or AES-CTR, which are widely supported. However, the root cause analysis shows that the real problems are in the SSH specification. We therefore suggest two changes to the specification, and a way to negotiate support for these changes without breaking backward compatibility.
@@
Resetting sequence numbers to zero when encryption keys are activated ensures that these sequence number manipulations during the handshake can no longer affect the secure channel. Resetting the sequence numbers is a major break in compatibility. To avoid connection failures due to only one peer resetting their sequence numbers, we suggest that an implementation signals the support for this countermeasure by including an identification string in the list of supported key exchange algorithms. The SSH extension negotiation mechanism is already employing this method. If and only if both peers signal support for this countermeasure, the sequence numbers will be reset.
Full Transcript Hash
The second countermeasure is to authenticate the full handshake transcript, as seen by the client and the server. This can detect any attempts of handshake manipulation by a MitM attacker, including sequence number manipulation through our techniques. It is not possible to simply extent the scope of the exchange hash, as the server signature is transmitted before the new keyset is taken into use. Therefore, any messages sent after the key exchange, but before NewKeys, can not be included. We suggest that the both peers send a MAC of the full transcript at the start of the secure channel, similarly to TLS FINISHED messages. Signaling support should be done as above.
@@
We suggest that SSH specifies “end-of-communication” messages to detect suffix truncation attacks. Also, AsyncSSH should be hardened to disallow unauthenticated, application-layer messages during the SSH handshake.
@@ 9 Conclusion
We introduced the novel sequence number manipulation and prefix truncation attacks for secure channels, which invalidate the INT-PST [15] security of SSH BPP for certain ciphers. We extended this vulnerability to real-world exploits like disabling SSH extension negotiation. This yields novel insights into the complex interplay between a practical security mechanism (sequence numbers) and abstract security notions (INT-PTXT vs. INT-CTXT vs. AEAD, [5]): Since implicit sequence numbers are not transmitted, they cannot be part of the ciphertext – therefore generic EtM modes fail, and INT-CTXT does not protect against prefix truncation. INT-PTXT and AEAD modes, on the other hand, do include them, either as part of the plaintext, or as associated data.
@@
Our close look at the Extension Negotiation mechanism reveals its design weaknesses: First, sending ExtInfo is optional even if both parties signal support for extension negotiation during the handshake. Second, SSH extension negotiation cannot be used to negotiate extensions affecting the SSH handshake itself, e.g., the countermeasures proposed in this paper. As a consequence, all extension negotiations should be done within the KexInit.
Although we suggest backward-compatible countermeasures to stop our attacks, we note that the security of the SSH protocol would benefit from a redesign from scratch. This redesign should be guided by all findings and insights from both practical and theoretical security analysis, in a similar manner as was done for TLS 1.3.
Appendix A Advanced Sequence Number Manipulation
Using that sequence numbers rollover to zero after $12^{32}-$ 1, the attacker can also attempt to decrement $\mathsf{C.Rcv}$ and $\mathsf{S.Rcv}$:
Technique RcvDecrease (5(b))
During the handshake, a MitM attacker can decrease $\mathsf{C.Rcv}$ (resp. $\mathsf{S.Rcv}$) by $N$, while not changing any other sequence number, by sending $2^{32}-N$ Ignore messages to the client (resp. server).
We can also combine these techniques to manipulate the $\mathsf{C.Snd}$ and $\mathsf{S.Snd}$ sequence numbers. For this, we require a message that generates a response message but is otherwise ignored. Conveniently, the SSH standard requires this for all messages with unrecognized message IDs [34, Sec. 11.4]. Let Unknown be a message with an unrecognized message ID.
Technique SndIncrease (5(c))
During the handshake, a MitM can increase $\mathsf{C.Snd}$ (resp. $\mathsf{S.Snd}$) by $N$ while not changing any other sequence number by sending $N$ Unknown and $2^{32}-N$ Ignore messages to the client (resp. server) and deleting all generated Unimplemented messages.
Technique SndDecrease (5(d))
During the handshake, a MitM can decrease $\mathsf{C.Snd}$ (resp. $\mathsf{S.Snd}$) by $N$ while not changing any other sequence number by sending $2^{32}-N$ Unknown and $N$ Ignore messages to the client (resp. server), and deleting all generated Unimplemented messages.
Here, the total data transfer required is $\approx 69\mathrm{GB}$ for SndIncrease and twice as much, or $\approx 138\mathrm{GB}$ for SndDecrease. Again, these techniques may fail on implementations that have timeouts or restrict the amount of data or number of messages exchanged during the handshake.
Evaluation

### v2

@@ 6.1 Rogue Extension Negotiation Attack
Figure 6: Rogue Extension Negotiation Attack on AsyncSSH: The MitM injects a malicious extension information message before the key exchange completes and deletes the server’s ExtInfo message to account for the change in sequence numbers. This attack relates to the generic extension downgrade attack in Section 5.2.
@@
Impact
The attacker can replace the content of the ExtInfo message. AsyncSSH clients support the server-sig-algs and global-requests-ok extensions. Hence, the attacker can try to downgrade the algorithm used for client authentication by restricting the value of server-sig-algs to a subset of those supported by the server.
Attack Description
@@ 6.2 Rogue Session Attack
Figure 7: Rogue Session Attack on AsyncSSH: The MitM injects a malicious authentication request before the handshake is complete and deletes the client’s ExtInfo message to account for the change in sequence numbers. By delaying the authentication request sent by the client, the MitM ensures that the malicious one is processed. Any additional authentication requests are silently ignored.
The rogue session attack targets any SSH client connecting to an AsyncSSH server, on which the attacker must have a shell account. The attack’s goal is to log the client into the attacker’s account without the client being able to detect this.
Impact
With a successful attack, the attacker can gain complete control over the remote end of the SSH session. The attacker receives all keyboard input by the user, completely controls the terminal output of the user’s session, can send and receive data to/from forwarded network ports, and can create signatures with a forwarded SSH Agent, if any. The result is a complete break of the confidentiality and integrity of the secure channel, providing a strong vector for a targeted phishing campaign against the user. For example, the attacker can display a password prompt and wait for the user to enter the password, elevating the attacker’s position to a MitM at the application layer and enabling impersonation attacks.
Attack Description
The messages exchanged during the attack are depicted in Figure 7. The attacker injects a chosen UserAuthRequest before the client’s NewKeys. This request must be a valid authentication request containing the credentials of the attacker. The attacker can use any authentication mechanism that does not require exchanging additional messages between client and server, such as password or publickey. Due to a state machine flaw, the AsyncSSH server accepts the unauthenticated UserAuthRequest message and defers it until the client has requested the authentication protocol.
To avoid a race condition between the UserAuthRequest sent by the client and the UserAuthRequest injected by the attacker, the attacker delays the client’s UserAuthRequest until after the server signals a successful authentication in response to the injected UserAuthRequest. The AsyncSSH server silently ignores any additional authentication request after a successful authentication.
@@
  • Alternatively, suppose the server sends an extra message before ServiceAccept. In that case, the attacker can delete that message after injecting an additional Unknown message to the client before NewKeys, triggering an Unimplemented response that is deleted. This increases both $\mathsf{C.Snd}$ and $\mathsf{C.Rcv}$, moving the send count deficit from the client to the server.
@@
We successfully evaluated the attack against AsyncSSH 2.13.2 as a server, connecting to AsyncSSH 2.13.2 and OpenSSH 9.4p1.
@@ 7 SSH Deployment Statistics
To estimate the impact of the prefix truncation attacks, we scan for the SSH servers preferring or supporting any affected encryption mode. Similarly, to estimate the impact of the extension downgrade attack, we scan for servers sending ExtInfo messages.
@@
For scanning, we used ZMap [16] and ZGrab2 [15] on port 22 of the entire IPv4 address space. The scan was performed over two days in early October 2023, totaling 15.164M SSH servers.
As ZGrab2 cannot capture SSH extensions, we performed a complementary scan at the end of June 2023, using a custom tool, on a subset of $2^{20}$ open ports. The scan covered a total of 830k servers. All data relating to the use of extension negotiation in SSH is sourced from this scan.
@@
In Table 2, we show the number of servers that prefer and support various encryption modes. A cipher is preferred if it is placed first in the list of supported algorithms.
We find that, by far, the most preferred encryption cipher is ChaCha20-Poly1305, with 57.64% listing this algorithm first. This is followed by AES-CTR (31.56%) and, with some distance, by AES-GCM (8.04%) and AES-CBC (1.56%).
@@
As non-AEAD ciphers must be combined with a MAC, we also evaluate which authenticated encryption modes the servers prefer and support. The numbers for the AEAD modes ChaCha20-Poly1305 (57.64%) and GCM (8.04%) are identical to those for encryption modes, as the MAC is already integrated. Preference for CTR modes is split between a majority for CTR-EaM (26.14%) and a minority for CTR-EtM (5.46%). Preference for CBC modes is mostly CBC-EaM (2.37%), while preference for CBC-EtM (0.09%) is marginal.
In summary, 63.2% of all servers prefer an authenticated encryption mode affected by our attacks.
Looking at the support for authenticated encryption modes vulnerable to our attacks, we find that 67.58% of all servers support ChaCha20-Poly1305, while 17.24% support CBC-EtM. In total, 71.6% support at least one affected mode.
AE Mode Preferred Supported
@@
Other 2k 0.01% - -
Unknown / No KexInit 36k 0.24% - -
@@
We also looked at SSH extensions offered by servers before user authentication; see Table 4. We can see that 76.81% of all servers send the server-sig-algs extensions to indicate support for better signature schemes for client public key authentication. Furthermore, 8.8% send the publickey-hostbound extension, improving security for authentication using SSH agent. Both extensions provide opportunities for downgrade attacks, as their absence can weaken the strength of the authentication.
@@
Table 4: SSH extensions offered by servers after the initial handshake, @openssh.com abbreviated to @. Extensions sent by servers upon successful client authentication are not included.
@@ 8 Suggested Countermeasures
As a stop-gap measure, the affected cipher modes can be turned off. Widely supported alternatives are AES-GCM or AES-CTR. However, the root cause analysis shows that the underlying issues lie in the SSH specification. We therefore suggest two changes to the specification.
@@
Resetting sequence numbers to zero when encryption keys are activated ensures that sequence number manipulations during the handshake can no longer affect the secure channel. Unfortunately, sequence number reset is a major break in compatibility. To avoid connection failures due to one-sided sequence number resets, we suggest that an implementation signals the support for this countermeasure by including an identification string in the list of supported key exchange algorithms. The SSH extension negotiation mechanism is already employing this method. If and only if both peers signal support for this countermeasure, the sequence numbers will be reset.
In response to our findings, OpenSSH implemented this behavior as part of their so-called “strict kex” countermeasure [36, Sec. 1.10]. In addition to resetting sequence numbers, “strict kex” mandates that unexpected or unknown messages during the initial key exchange must lead to the connection’s termination. An unexpected message in this context is any message that is not strictly required for key exchange. “strict kex” has since been adopted by various vendors to ensure interoperability between SSH implementations.
Full Transcript MAC
Authenticating the full handshake transcript, as seen by the client and server, can detect attempts of handshake manipulation by a MitM attacker, including sequence number manipulation through our techniques. It is impossible to extend the scope of the existing exchange hash, as the server signature is transmitted before the new keys are taken into use. Therefore, any messages sent after the key exchange but before NewKeys cannot be included. We suggest that both peers send a MAC authenticating the entire transcript at the start of the channel, similar to TLS Finished messages. Signaling support should be done as above. However, the transcript must be carefully canonicalized. While client and server messages are sequential, they can interleave asynchronously, leading to transcript variations. Also, the protocol must be extended to define the algorithm, encoding, and position of the transcript MAC. Thus, securing the handshake is more complex than resetting the sequence number.
Relationship to Formal Proofs
Both countermeasures have a common goal: Align the SSH standard with expectations for stateful encryption schemes from formal models for the BPP presented in [4, 1]. A sequence number reset achieves this directly by initializing the sequence numbers to zero, as in the models. On the other hand, verifying the full transcript hash forces the sequence number in the stateful encryption and decryption methods to be synchronized by the sender and receiver. Although the sequence numbers are then not initialized to zero, each pair is nevertheless initialized to a common value out of the attacker’s control. The existing models could then be adjusted in the following way: If $T_{C},T_{S}\in\{0,1\}^{*}$ are the (canonicalized) transcripts of the SSH handshake as seen by the client and the server, and $M_{CS},M_{SC}:\{0,1\}^{*}\mapsto\mathbb{N}$ are functions counting the messages from the client to the server and vice versa in a transcript, then sequence numbers in the stateful encryption and decryption modes are initialized to:
$\mathsf{C.Snd}$ $=M_{CS}(T_{C}),$ $\mathsf{C.Rcv}$ $=M_{SC}(T_{C}),$
$\mathsf{S.Snd}$ $=M_{SC}(T_{S}),$ $\mathsf{S.Rcv}$ $=M_{CS}(T_{S}).$
Authenticating the transcript then ensures that $T_{C}=T_{S}$, and thus $\mathsf{C.Snd}=\mathsf{S.Rcv}$ and $\mathsf{C.Rcv}=\mathsf{S.Snd}$, before the first messages in the secure channel are encrypted or decrypted. Authenticating the handshake transcript has the added benefit that the handshake could be analyzed in a “matching conversations”-based security model [23, 24].
@@
We suggest that SSH specifies “end-of-communication” messages to detect suffix truncation attacks. Also, AsyncSSH should be hardened to disallow unauthenticated, application-layer messages during the SSH handshake. In response to our findings, the state machine of AsyncSSH was improved in version 2.14.1 to mitigate our attacks.
9 Future Work
Formally, SSH BPP security was modeled as stateful decryption [4, 40, 1]. Implicitly, this state was associated with SSH sequence numbers, and it was assumed that an adversary could not manipulate this state. These models can be extended in two directions: (1) Include a broader definition of state. By including chained IVs, key stream state, and GCM invocation counters, these models can be used to show why certain cipher modes resist our attacks and that they indeed achieve INT-PST security. (2) Introduce a novel adversarial query, ModifyState, to model the attacks described here.
Our attack combines weaknesses in the SSH handshake with weaknesses in the encrypted channel. Earlier work analyzed these separately, leading to small models. To find our attack automatically, models of SSH for computer-aided proofs could (1) model the handshake as well as the BPP together, (2) keep track of sequence numbers in the BPP, including the handshake, which requires modeling integer numbers that can overflow as the internal state, (3) model seemingly unimportant messages like Ignore, and (4) consider each encryption mode separately. The properties to verify should include strong security notions such as INT-aPTXT [17].
Applying state learning to implementations also has the potential to find our attacks automatically in the future, although it suffers from a combinatorial explosion in the number of messages (see Section 5.1 in [28]). Messages like Ignore and ExtInfo need to be included in the alphabet to find our attacks, and an active MitM attacker has to be considered.
@@ 10 Conclusion
We introduced novel sequence number manipulation and prefix truncation attacks for secure channels, which invalidate the INT-aPTXT [17] security of SSH BPP for certain ciphers. We extended these vulnerabilities to real-world exploits like disabling SSH extension negotiation. This yields novel insights into the complex interplay between a practical security mechanism (sequence numbers) and abstract security notions (INT-PTXT vs. INT-CTXT, [6]).
@@
Our close look at the extension negotiation mechanism reveals its design weaknesses: First, sending ExtInfo is optional even if both parties signal support for extension negotiation during the handshake. Second, ExtInfo cannot be used to change the SSH handshake itself, e.g., to implement the countermeasures proposed in this paper. However, it outperforms extension negotiation within the KexInit in aspects of privacy as protocol extension can be negotiated securely, i.e., privately, similar to encrypted extensions in TLS 1.3. As a consequence, extension negotiation within the KexInit should be strictly limited to extensions affecting the SSH handshake. Protocol extensions affecting the user authentication or application layer should be negotiated through the extension negotiation mechanism.
Although we suggest backward-compatible countermeasures to stop our attacks, the security of the SSH protocol could benefit from a redesign from scratch. The redesign process could be inspired by that of TLS 1.3, which brought implementers together with experts in protocol analysis and formal verification [3]. This could simplify the protocol while preserving and/or achieving desired security notions for SSH, which may differ from those of TLS. For example, while the privacy of client authentication and extension negotiation are relatively new features for TLS, they are already present in SSH and should thus be preserved in a redesign.
Acknowledgements
Fabian Bäumer was supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) project “Industrie 4.0 Recht-Testbed” (13I40V002C). Marcus Brinkmann was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972.
Appendix A Artifact Appendix
A.1 Abstract
This document describes the artifacts to the USENIX Security ’24 Publication Terrapin Attack: Breaking SSH Channel Integrity By Sequence Number Manipulation.
Using these instructions, the evaluations of Sequence Number Manipulation (Sect. 4.1), Extension Downgrade Attack (Sect. 5.2), Rogue Extension Attack (Sect. 6.1) and Rogue Session Attack (Sect. 6.2) can be reproduced.
Also, the aggregation scripts for the internet scans are available and can be tested on a small subset of the samples.
A.2 Description & Requirements
A.2.1 Security, privacy, and ethical concerns
The configuration uses the host network to allow (optional) monitoring of the attack using Wireshark or other network packet analysis tools on the loopback interface. During the runtime of the evaluation, this makes the tested SSH server and proof of concept (PoC) available to all systems with access to the local network (TCP bind to 0.0.0.0, ports 2200 and 2201). Reviewers should take care to isolate the test system from the internet, for example using a firewall.
A.2.2 How to access
A.2.3 Hardware dependencies
None.
A.2.4 Software dependencies
  • Linux or MacOS. [^6] Windows WSL might work but is untested and not supported. No specific distribution or version is required. We used Manjaro (rolling release in March 2024) and MacOS 14.4 (Sonoma).
  • Bash shell interpreter (typically included in the above). No specific version is required. We used bash 5.2.26 and 3.2.57.
  • Docker Engine or Docker Desktop. While Docker Engine suffices and is typically included in Linux distributions, Docker Desktop is a separate install on MacOS. No specific version is required. We used Docker Engine 25.0.3 and Docker Desktop 4.28.0.
  • Wireshark (optional), for network packet analysis.
A.2.5 Benchmarks
None.
A.3 Set-up
All required Docker images are built on demand when the evaluation scripts are executed, so no setup is required.
The TCP ports 2200 and 2201 should be free and available. This is the case by default on many systems. Some systems might require a configuration of the firewall to allow the test servers to bind 0.0.0.0 on these ports. On some systems, the firewall will show a pop-up dialog when the first server starts up, requiring manual confirmation.
A.3.1 Installation
Linux:
Install the Docker engine under a supported Linux distribution by following the instructions available at https://docs.docker.com/engine/install/.
MacOS:
Install Docker Desktop available at https://www.docker.com/products/docker-desktop/.
A.3.2 Basic Test
The following scripts build all required Docker images and can be used as a basic functionality test. It will also be called by all evaluation scripts, so this step is optional.
1\$ impl/build.sh
2[+] Building 2.0s (15/15) FINISHED
3[...]
4=> => naming to docker.io/terrapin-artifacts/openssh-server:9.4p1
5[...]
6\$ pocs/build.sh
7[...]
The output shows the progress on downloading base images and building the evaluation images. If there is no output, all docker images are already built.
A.4 Evaluation workflow
A.4.1 Major Claims
We evaluated our attacks against several clients using an OpenSSH 9.5p1 (C1, C2) or AsyncSSH 2.13.2 (C3, C4) server. For an overview of the expected outcomes, see also Table 5.
Attack PuTTY 0.79 OpenSSH 9.4p1 OpenSSH 9.5p1 Dropbear 2022.83 AsyncSSH 2.13.2 libssh 0.10.5
C1 RcvIncrease -
C1 RcvDecrease - R T T
C1 SndIncrease - R U T T
C1 SndDecrease - R U T T
C2 ChaCha-Poly - - -
C2 CBC-EtM
C2 - Unknown 0.0300 0.0003 0.0003 - - -
C2 - Ping 0.8383 - 0.0074 - - -
C3 Rogue Extension - - - - -
C4 Rogue Session - - - - -
    Not evaluated.
  • Attack succeeds.
  • R
    Client terminates the connection (rollover).
  • T
    Client terminates the connection (timeout).
  • U
    Client terminates the connection (Unknown message).
Table 5: Expected outcomes for attacks against clients
(C1):